# Imaginaries, products and the adele ring

### v2

@@ 2.6. Expansions of $T^{boole}$: atomic, atomless and atomic-by-atomless
    $T^{boole}_{{atomless}}$ asserting in addition to $T^{boole}$ that $B$ is atomless. Again we will shorten the notation to $T_{{atomless}}$
@@
    Let $L_{mixed}$ be $L^{boole}$ expanded by a predicate $J\subset B$, and a sort $\bar{B}$ for $B/J$; and let $T_{mixed}=T^{boole}+{ABAI}_{\infty}$. In other words $T_{mixed}$ asserts that $J$ is an ideal of $B$ including all atoms, and that $B/J$ is atomless.
    We adjoin to $T_{mixed}$ two copies of each sorts of L{L} - one to be interpreted as $B$-valued, the other as $\bar{B}$}-valued, i.e. the reduced (Boolean) product structure; along with the sorts $B$ and $\bar{B}$} (they can be viewed as a special case, if L{L} is taken to have a two-element sort with elements $0,1$.)
By Theorem 2.3 and Lemma 2.5, all three admit quantifier-elimination. If $T$ is complete, then so is $T_{atomless}$, while $T_{mixed}$ and $T_{at}$ become complete upon specifying that there are infinitely many atoms.
Recall that $M/J$ is defined as $M/\sim_{J}$, where $\sim_{J}$ is the equivalence relation defined by $x\sim_{J}y$ iff $x\neq y]\in J$. We’ll sometimes write "mod JJ" as a shorthand for "mod $\sim_{J}$".
@@
Let $M\models T_{mixed}$. Define
$\phi]:(M/J)^{n}\to B/J$
by
$\phi](a\mod J)=[\phi](a)\mod J$
Then $\bar{M}=M/J$ is a model of $T_{atomless}$; it is embedded and stably embedded in $M$.
@@
We will denote the quotient maps $M\to\bar{M}$ and $B\to\bar{B}$ by $\pi$ . We will prove quantifier-elimination for $(M,\bar{M},\pi,B,\bar{B},\pi,[\phi]:M^{n}\to B,[\phi]:\bar{M}^{n}\to\bar{B})$ (where $\phi$ ranges over all formulas of $L$.) From this it is clear that $M/J$ is stably embedded, and that the induced structure is the specified one.
It is clear that $\bar{M}$ is a model of $T_{atomless}$, and so we already have quantifier elimination for $(\bar{M},\bar{B},[\phi]:\bar{M}^{n}\to\bar{B})$.
It remains to eliminate one existential quantifier over $M$. Thus we must decide whether there exists $x\in M$ with $\pi(x)$ specified as a given element $\mathbf{a}$} of $M/J$, and with $\phi_{i}(x)]=b_{i}$ for $i=1,\ldots,n$. Here we may assume $\phi_{1},\ldots,\phi_{n}$ are a partition of unity in $T$. A necessary condition is that $\phi_{i}({\mathbf{a}})]=\pi(b_{i})$ for each $i$. Conversely assuming this condition holds, lifting $\mathbf{a}$} arbitrarily to an element $a^{\prime}$ of $M$, we have $\pi(a^{\prime})={\mathbf{a}}$; and for some element $e$ of $J$, $\phi_{i}(a^{\prime})]=b_{i}(\mod e)$ for each $i$. Using Axiom 2.2 (4), find $a^{\prime\prime}$ such that $\phi_{i}(a^{\prime\prime})]\cap e=b_{i}\cap e$. Next glue $a^{\prime},a^{\prime\prime}$ together, i.e. find $a$ with $a=a^{\prime\prime}]=e$ and $a=a^{\prime}]=1-e$. Then $\pi(a)={\mathbf{a}}$ and $\phi_{i}(a)]=b_{i}$, as required. ∎
@@
Note that $B$ is interpretable in the original sorts of $M$, provided ${T}\models(\exists x)(\exists y)(x\neq y)$ for variables $x,y$ of some sort $S$ of $L$, the language of $T$. In this case an element $a\in{\mathop{\operatorname{\rm P}}}(I)$ can be coded as the equalizer of a pair of elements $c,d$ in $S$, i.e. $a=\{i:c(i)=d(i)\}$. If we assume further that $L$ has two constant symbols for elements of $S$, then $B$ is not only interpretable but definable. In any case we will view $B$ as a sort in the language of the product.
The Boolean algebra of $\prod_{i\in I}M_{i}$, namely ${\mathop{\operatorname{\rm P}}}(I)$, is atomic. If $I_{at}$ is the ideal generated by the atoms in $B(M)$, then the reduced product $\prod_{i\in I}M_{i}/I_{at}$ has atomless Boolean algebra, and becomes a model of $T_{{atomless}}$. Hence if $I$ is infinite, with $J:=I_{at}$, $M$ becomes a model of $T_{mixed}$.
@@
    Let $x$ be a variable on the main sort $S$ of $T_{atomless}$; then the quantifier $(\exists^{\infty}x)$ is definable. Moreover for $a\subset M^{k}$, $M\models T_{atomless}$, $acl(a)\cap S=dcl(a)\cap S$.
@@
where $\Theta$ is a formula of $(B,J)$ and $\phi_{1},\ldots,\phi_{K}$ formulas of $L$, forming a partition of unity. Replace each $\phi_{k}(x)$ by the two formulas:
@@
To achieve the ’moreover’ is easy in case $\phi_{k_{2}}$ has >n!>n! solutions in each $M_{i}$, $i\in I_{1}$. It suffices in this case, having picked $c_{i}$, to pick $d_{i}$ to differ from any of the at most $n!$ permutations of $d_{i}$. Since we only know that $\phi_{k_{2}}$ has at least two solutions in each $M_{i},i\in I_{1}$, we use a trick. Let $l$ be an integer with $2^{l}>n!$, and divide $I_{1}$ into blocks of $l$ elements; so that $I_{1}=[l]\times I_{1}^{\prime}$. We will define $c_{i},d_{i}$ not one index $i$ at a time but in blocks of $l$ indices. For $i^{\prime}\in I_{1}^{\prime}$, let $M_{i^{\prime}}=\prod_{j\in[l]}M_{(j,i^{\prime})}$ be the product of the $l$ elements of the $i^{\prime}$’th block. Let $I^{\prime}=(I\setminus I_{1})\cup I_{1}^{\prime}$ and $i^{\prime}=i$ for $i\notin I_{1}$. Let $\phi_{k_{2}}^{\prime}$ be the product over the block of $\phi_{k_{2}}$. Then $\prod_{i\in I}M_{i}=\prod_{i^{\prime}\in I^{\prime}}M_{i^{\prime}}$. But now $\phi_{k_{2}}^{\prime}$ has at least $2^{l}>n!$ solutions in each $M_{i^{\prime}}$, so we can choose $c_{i^{\prime}},d_{i^{\prime}}$ so that no $Sym(n)$-permutation of $c_{i^{\prime}}$ equals $d_{i^{\prime}}$. This achieves also the ‘moreover’.
(4) Similar to the first part of (3). We can also view $S$ as a sort of $T_{mixed}$ (the main sort modulo $J$.) We have an $a$-definable set $D=D_{a}$ and we wish to find conditions on $a$ for finiteness of $D$. We may again write $y\in D$ iff $\Theta([\phi_{1}(y)],\cdots,[\phi_{K}(y)])$, where now the $\phi_{i}$ depend upon the parameter $a$. The analysis of (3) shows uniformly in $a$ that either $D_{a}$ is infinite, or $D_{a}$ decomposes explicitly into a bounded finite number of $a$-definable sets with a single element. Hence algebraic and definable closures coincide on $S$. $T_{mixed}$
Recall that a theory $T$ admits weak elimination of imaginaries iff for any imaginary element $e$ in any model of $T$, we have $e\in dcl(e_{1},\ldots,e_{k})$ for some ‘real’ $e_{1},\ldots,e_{k}\in acl(e)$. Equivalently, any imaginary $e$ is equi-definable with a finite set of tuples real elements.
See e.g. [5], 2.3 for this notion as well as stable embeddedness.
Our main goal will be the proof of weak elimination of imaginaries for $T_{at}$, $T_{atomless}$, and $T_{mixed}$, given elimination of imaginaries for $T$. This will be achieved in Theorems 3.1, 4.1, and 5.1, respectively. The remark below explains why assuming full EI for $T$ is necessary.
@@
For instance, let $4$ denote a four-element pure set; let $4\choose 2$} be the set of $2$-element subsets of $4$; then $4^{\mathbb{N}}$ has $4\choose 2$$^{\mathbb{N}}$ as an imaginary sort; let $e\in{4\choose 2}^{\mathbb{N}}$, viewed as an imaginary of $4^{\mathbb{N}}$; then no element of $4^{\mathbb{N}}$ is algebraic over $e$, so $4^{\mathbb{N}}$ does not admit weak elimination of imaginaries. Nor does the theory with $\clubsuit$ obtained by adding an ideal element, as in Remark 1.3. (We leave these statements to the reader, noting that if $G$ acts on $X$ without fixed points, then $G^{\mathbb{N}}$ acts on $X^{\mathbb{N}}$ without finite orbits. Apply this to $G=Aut(4;\{1,2\})$ acting on $4$.)
@@ 3. Imaginaries in infinite products
Following some preliminaries, we will prove the theorem for the very special case of the theory $T=SET_{n}$ of an $n+1$-element set, with a single distinguished point (Lemma 3.13.) Here we will use liaison groups in a somewhat new setting, and make use of results of descriptive set theory. From this the case of the theory of a finite set with arbitrary structure follows immediately (Corollary 3.14). We will then prove Theorem 3.1 in general, by an argument reducing to the case of products of finite structures.
@@
$\clubsuit$] Assume $I$ is finite, and $M_{i}$ has weak EI for $i\in I$. Consider $M=\prod_{i\in I}M_{i}$ as a structure as in 2.8. Then $M$ admits weak EI.
@@
Note that the structure of 2.8 includes, in this case, the finite Boolean algebra ${\mathop{\operatorname{\rm P}}}(I)$. We have in general:
@@
Let $Z$ be a definable subset of $M^{n}$ , with canonical parameter $e\in M^{eq}$; we will find a finite subset $W_{i}$ of $M_{i}$ such that $e$ is interdefinable with a code for a certain subset $W$ of $W_{1}\times W_{2}$. We note first that $Z$ is canonically a finite union of some rectangles $X_{i}\times Y_{j}$ with $X_{1},\cdots,X_{k}$ definable in $M_{1}$ and $Y_{1},\ldots,Y_{l}$ definable in $M_{2}$. Namely, let $\mathcal{Y}_{2}=\{Z(a):a\in M_{1}\}$; where $Z(a)=\{y:(a,y)\in Z\}$. Then $\mathcal{Y}_{2}$ is a finite set of $M_{2}$-definable subsets of $M_{2}$. (To see this, it suffices to saturate $(M_{1},M_{2},M_{1}\times M_{2})$ and then note that $Y_{2}$ is bounded. This is because an automorphism of $M_{1}$ extends to an automorphism of $M_{1}\times M_{2}$ that fixes $M_{2}$. So $Z(a)=Z(b)$ if $tp_{M_{1}}(a)=tp_{M_{1}}(b)$.)

### v3

@@ 2.6. Expansions of $T^{boole}$: atomic, atomless and atomic-by-atomless
    $T^{boole}_{atomless}$ asserting in addition to $T^{boole}$ that $B$ is atomless.
@@
    Let $L_{mixed,0}$ be $L^{boole}$ expanded by a predicate $J\subset B$. $T_{mixed,0}$ includes $T^{boole}$ and asserts in addition that $B$ is atomic with infinitely many atoms, that $J$ is a proper ideal of $B$ with is atomless.
    Note that it follows that $J$ includes all atoms of $B$.
    $L_{mixed}$ will be an inessential expansion of $L_{mixed,0}$. $L_{mixed,0}$ includes a sort $S_{B}$ for each sort $S$ of $L$. We define $L_{mixed}$ by adjoining to $L_{mixed,0}$ a sort $\bar{B}$ along with a map $\pi_{B}:B\to\bar{B}$; and for each sort $S$ of $L$, a sort $S_{\bar{B}}$ and a map $\pi_{S}:S_{B}\to S_{\bar{B}}$.
    Let $\sim_{J}$ be the equivalence relation defined by $x\sim_{J}y$ iff $x\neq y]\in J$.
    $T_{mixed}$ extends $T_{mixed,0}$ by the assertions that the maps $\pi_{B}$ and $\pi_{S}$ are surjective with kernel given by $\sim_{J}$.
We will put together all the $\pi_{S}$ and denote them by $\pi$.
We’ll often write $\mod\,J$" as a shorthand for $\mod\,\sim_{J}$", and in particular $M/J$ for $M/\sim_{J}$.
Remark 2.7.
If L{L} has a two-element sort $\mathsf{\bf{2}}$ with elements $0,1$, the Boolean algebra $B$ of (1,2), and the algebras $B$ and $\bar{B}$} of (3), do not require separate treatment; they can be identified with the Boolean interpretation of $\mathsf{\bf{2}}$.
By Theorem 2.3 and Lemma 2.5, all three admit quantifier-elimination. If $T$ is complete, then so is $T_{atomless}^{boole}$, while $T_{mixed}$ and $T_{at}$ become complete upon specifying that there are infinitely many atoms.
For each $L$-formula $\phi$, the function $\phi]:M^{n}\to B$ induces a function $(M/J)^{n}\to B/J$, that we will denote $\bar{\phi}]$. The defining equation is:
$\bar{\phi}](a\mod J)=[\phi](a)\mod J$
We will denote both quotient maps $M\to\bar{M}$ and $B\to\bar{B}$ by $\pi$; specifying the domain and range when needed, but the domains are in different sorts so there is no danger of confusion.
@@
Let $M\models T_{mixed,0}$, and let
$\widetilde{M}=(M,\bar{M},B,\bar{B},\pi:B\to\bar{B},\pi:M\to\bar{M},[\phi]:M^{n}\to B,[\bar{\phi}]:\bar{M}^{n}\to\bar{B})$
(where $\phi$ ranges over all formulas of $L$) be the corresponding model of $T_{mixed}$. Then:
  1. (1)
    $\bar{M}=M/J$ is a model of $T_{atomless}^{boole}$.
  2. (2)
    Quantifier-elimination holds for $\widetilde{M}$.
  3. (3)
    $M/J$ is embedded and stably embedded in $\widetilde{M}$.
@@
It is clear that (1) holds, $\bar{M}$ is a model of $T_{atomless}^{boole}$. Thus we already have:
(2.1) Quantifier elimination for $(\bar{M},\bar{B},[\bar{\phi}]:\bar{M}^{n}\to\bar{B})$.
We now prove (2). In view of (2.1), it remains to eliminate quantifiers over the sorts of $M$; this reduces to eliminating one existential quantifier over $M$. Consider a quantifier-free formula $\theta(x,u)$ with $x$ ranging over $M$. We need to find a quantifier-free formula equivalent to $(\exists x)\theta(x,u)$. Equivalently, if we add constant symbols $d$ to the language matching $u$, we must find quantifier-free formula equivalent to $(\exists x)\theta(x,d)$. We thus supress the reference to $u$ in the notation.
The variable $x$ can only appear via $\pi(x)$ and various $\phi(x)]$; i.e. $\theta(x)$ is equivalent to
$\theta^{\prime}([\phi_{1}(x)],\ldots,[\phi_{n}(x)],\pi(x))$
where $x$ does not occur in $\theta^{\prime}$. So we can rewrite $\theta$ as:
$(\exists b_{1}\in B)\ldots(\exists b_{n}\in B)(\exists\bar{a}\in M/J)\bigwedge_{i=1}^{n}[\phi_{i}(x)]=b_{i}\wedge\pi(x)=\bar{a}\wedge\theta^{\prime}(b_{1},\ldots,b_{n},\bar{a})$
Let $\psi(b_{1},\ldots,b_{n},\bar{a}):=(\exists a)(\pi(a)=\bar{a}\wedge\bigwedge_{i=1}^{n}[\phi_{i}(a)]=b_{i}$). Then $(\exists x)\theta$ is equivalent to
$(\exists b_{1}\in B)\ldots(\exists b_{n}\in B)(\exists\bar{a}\in M/J)\psi(b_{1},\ldots,b_{n},\bar{a})\wedge\theta^{\prime}(b_{1},\ldots,b_{n},\bar{a}).$
We will show that $\psi$ is equivalent to a quantifier-free formula; given this, the quantifiers on $b_{1},\ldots,b_{n},\bar{a}$ can be eliminated using (i).
Thus we must decide (on the basis of a quantifier-free formula), given an element $\bar{a}$ of $M/J$ and elements $b_{1},\ldots,b_{n}\in B$, whether there exists $x\in M$ with $\pi(x)=\bar{a}$ and $\phi_{i}(x)]=b_{i}$ for $i=1,\ldots,n$. Here we may assume $\phi_{1},\ldots,\phi_{n}$ are a partition of unity in $T$.
A necessary condition is that $\bar{\phi}_{i}(\bar{a})]=\pi(b_{i})$ for each $i$. Conversely assuming this condition holds, lifting $\bar{a}$ arbitrarily to an element $a^{\prime}$ of $M$, we have $\pi(a^{\prime})=\bar{a}$; and for some element $e$ of $J$, $\phi_{i}(a^{\prime})]=b_{i}(\mod e)$ for each $i$. Using Axiom 2.2 (4), find $a^{\prime\prime}$ such that $\phi_{i}(a^{\prime\prime})]\cap e=b_{i}\cap e$. Next glue $a^{\prime},a^{\prime\prime}$ together, i.e. find $a$ with $a=a^{\prime\prime}]=e$ and $a=a^{\prime}]=1-e$. Then $\pi(a)=\bar{a}$ and $\phi_{i}(a)]=b_{i}$, as required.
(3) From the nature of the quantifier-elimination (2) it is clear that $M/J$ is stably embedded, and that the induced structure is the specified one.
@@
Note that $B$ is interpretable in the original sorts of $M$, provided ${T}\models(\exists x)(\exists y)(x\neq y)$ for variables $x,y$ of some sort $S$ of $L$, the language of $T$. In this case an element $a\in{\mathop{\operatorname{\rm P}}}(I)$ can be coded as the equalizer of a pair of elements $c,d$ in $S$, i.e. $a=\{i\in I:c(i)=d(i)\}$. If we assume further that $L$ has two constant symbols for elements of $S$, then $B$ is not only interpretable but definable. In any case we will view $B$ as a sort in the language of the product.
The Boolean algebra of $\prod_{i\in I}M_{i}$, namely ${\mathop{\operatorname{\rm P}}}(I)$, is atomic. If $I_{at}$ is the ideal generated by the atoms in $B(M)$, then the reduced product $\prod_{i\in I}M_{i}/I_{at}$ has atomless Boolean algebra, and becomes a model of $T_{{atomless}^{boole}}$. Hence if $I$ is infinite, with $J:=I_{at}$, $M$ becomes a model of $T_{mixed}$.
@@
    Let $x$ be a variable on the main sort $S$ of $T_{atomless}^{boole}$; then the quantifier $(\exists^{\infty}x)$ is definable. Moreover for $a\subset M^{k}$, $M\models T_{atomless}^{boole}$, $acl(a)\cap S=dcl(a)\cap S$.
@@
where $\Theta$ is a formula of $(B,J)$ and $\phi_{1},\ldots,\phi_{K}$ formulas of $L$, forming a partition of unity. Replace each $\phi_{k}(x)$ by the two formulas
@@
For the ‘moreover’ part of the statement we argue as follows. In case $\phi_{k_{2}}$ has >n!>n! solutions in each $M_{i}$, $i\in I_{1}$, it suffices, having picked $c_{i}$, to pick $d_{i}$ to differ from any of the at most $n!$ permutations of $d_{i}$. Since we only know that $\phi_{k_{2}}$ has at least two solutions in each $M_{i},i\in I_{1}$, we use a trick. Let $l$ be an integer with $2^{l}>n!$, and divide $I_{1}$ into blocks of $l$ elements; so that $I_{1}=[l]\times I_{1}^{\prime}$. We will define $c_{i},d_{i}$ not one index $i$ at a time but in blocks of $l$ indices. For $i^{\prime}\in I_{1}^{\prime}$, let $M_{i^{\prime}}=\prod_{j\in[l]}M_{(j,i^{\prime})}$ be the product of the $l$ elements of the $i^{\prime}$’th block. Let $I^{\prime}=(I\setminus I_{1})\cup I_{1}^{\prime}$ and $i^{\prime}=i$ for $i\notin I_{1}$. Let $\phi_{k_{2}}^{\prime}$ be the product over the block of $\phi_{k_{2}}$. Then $\prod_{i\in I}M_{i}=\prod_{i^{\prime}\in I^{\prime}}M_{i^{\prime}}$. But now $\phi_{k_{2}}^{\prime}$ has at least $2^{l}>n!$ solutions in each $M_{i^{\prime}}$, so we can choose $c_{i^{\prime}},d_{i^{\prime}}$ so that no $Sym(n)$-permutation of $c_{i^{\prime}}$ equals $d_{i^{\prime}}$. This concludes the proof of the ‘moreover’ part.
(4) Similar to the first part of (3). We can also view $S$ as a sort of $T_{mixed}$ (the main sort modulo $J$.) We have an $a$-definable set $D=D_{a}$ and we wish to find conditions on $a$ for finiteness of $D$. We may again write $y\in D$ iff $\Theta([\phi_{1}(y)],\cdots,[\phi_{K}(y)])$, where now the $\phi_{i}$ depend upon the parameter $a$. The analysis of (3) shows uniformly in $a$ that either $D_{a}$ is infinite, or $D_{a}$ decomposes explicitly into a bounded finite number of $a$-definable sets with a single element. Hence algebraic and definable closures coincide on $S$.
Recall that a theory $T$ admits weak elimination of imaginaries iff for any imaginary element $e$ in any model of $T$, we have $e\in dcl(e_{1},\ldots,e_{k})$ for some ‘real’ $e_{1},\ldots,e_{k}\in acl(e)$. Equivalently, any imaginary $e$ is equi-definable with a finite set of tuples real elements. For more details on stable embeddedness see e.g. [6] (2.3); for weak elimination of imaginaries see e.g. [4]. Our main goal will be the proof of weak elimination of imaginaries for $T_{at}$, $T_{atomless}^{boole}$, and $T_{mixed}$, given elimination of imaginaries for $T$. This will be achieved in Theorems 3.1, 4.1, and 5.1, respectively. The remark below explains why assuming full EI for $T$ is necessary.
@@
For instance, let $4$ denote a four-element pure set and $4\choose 2$} the set of $2$-element subsets of $4$. Then $4^{\mathbb{N}}$ has $4\choose 2$$^{\mathbb{N}}$ as an imaginary sort. Let $e\in{4\choose 2}^{\mathbb{N}}$, viewed as an imaginary of $4^{\mathbb{N}}$. Then no element of $4^{\mathbb{N}}$ is algebraic over $e$, so $4^{\mathbb{N}}$ does not admit weak elimination of imaginaries. Nor does the theory with $\clubsuit$ obtained by adding an ideal element, as in Remark 1.3. (We leave these statements to the reader, noting that if $G$ acts on $X$ without fixed points, then $G^{\mathbb{N}}$ acts on $X^{\mathbb{N}}$ without finite orbits. Apply this to $G=Aut(4;\{1,2\})$ acting on $4$.)
@@ 3. Imaginaries in infinite products
Following some preliminaries, we will prove the theorem for the very special case of the theory $T=SET_{n}$ of an $n$-element set, with a single distinguished point (Lemma 3.13.) Here we will use liaison groups in a somewhat new setting, and make use of results of descriptive set theory. We will then prove Theorem 3.1 in general, by an argument reducing to enrichments of $(SET_{n})_{at}$.
@@
$\clubsuit$] Assume $I$ is finite, and $M_{i}$ has weak EI for $i\in I$. Consider $M=\prod_{i\in I}M_{i}$ as a structure as in 2.9. Then $M$ admits weak EI.
@@
Note that the structure of Example 2.9 includes, in this case, the finite Boolean algebra ${\mathop{\operatorname{\rm P}}}(I)$. We have in general:
@@
Let $Z$ be a definable subset of $M^{n}$ , with canonical parameter $e\in M^{eq}$; we will find a finite subset $W_{i}$ of $M_{i}$ such that $e$ is interdefinable with a code for a certain subset $W$ of $W_{1}\times W_{2}$. We note first that $Z$ is canonically a finite union of some rectangles $X_{i}\times Y_{j}$ with $X_{1},\cdots,X_{k}$ definable in $M_{1}$ and $Y_{1},\ldots,Y_{l}$ definable in $M_{2}$. Namely, let $\mathcal{Y}_{2}=\{Z(a):a\in M_{1}\}$; where $Z(a)=\{y:(a,y)\in Z\}$. Then $\mathcal{Y}_{2}$ is a finite set of $M_{2}$-definable subsets of $M_{2}$. (To see this, it suffices to saturate $(M_{1},M_{2},M_{1}\times M_{2})$ and then note that $\mathcal{Y}_{2}$ is bounded. This is because an automorphism of $M_{1}$ extends to an automorphism of $M_{1}\times M_{2}$ that fixes $M_{2}$. So $Z(a)=Z(b)$ if $tp_{M_{1}}(a)=tp_{M_{1}}(b)$).